OpenStudy (anonymous):

Can someone please help me and check my math test. It's only 10 questions. Btw, I do give medals. :)

OpenStudy (anonymous):

What is the value of x in the equation −6 + x = −3? −9 −3 3 (my choice) 9 Question 2 (Multiple Choice Worth 5 points) (07.05 MC) Which of these is a simplified form of the equation 7y + 8 = 9 + 3y + 2y? 7y = 6 2y = 1 (my choice) 12y = 17 5y = 17